solve the system by the elimination method 2x+3y=0 6x+11=6y
Answer by Alan3354(69443)

solve the system by the elimination method 2x+3y=0 6x+11=6y
--------------
2x+3y=0
6x+11=6y --> 6x-6y = 11
-----------
Multiply the 1st eqn by 2, then add the equations.